1.    10 fair die are rolled. Find the expected value of the sum of the faces showing.

2.    An urn contains 2 red chips and 3 white chips.

a.    Two are drawn out at random without replacement. Let X denote the number of red chips in the sample. Calculate Var(X).

b.    Calculate Var(X) if the sampling is done with replacement.


1
Expert's answer
2021-10-29T03:02:13-0400

1. For onefair dice


"E(X)=\dfrac{1}{6}(1)+\dfrac{1}{6}(2)+\dfrac{1}{6}(3)+\dfrac{1}{6}(1)+\dfrac{1}{6}(2)+\dfrac{1}{6}(3)"

"=3.5"

Expected value is linear, meaning that the expected value of the sum of random variables is the sum of their expected values. This means that if you roll "n"  dice, the expected value of the sum of the faces is "n" times the expected value of a single face, or "3.5n."

For 10 fair die

"E(Y)=10(3.5)=35"

The expected value of the sum of the faces showing is "35."


2.

a.  Without replacement.


"P(X=0)=\dfrac{\dbinom{2}{0}\dbinom{3}{2}}{\dbinom{2+3}{2}}=\dfrac{1(3)}{10}=\dfrac{3}{10}"

"P(X=1)=\dfrac{\dbinom{2}{1}\dbinom{3}{1}}{\dbinom{2+3}{2}}=\dfrac{2(3)}{10}=\dfrac{6}{10}"

"P(X=2)=\dfrac{\dbinom{2}{2}\dbinom{3}{0}}{\dbinom{2+3}{2}}=\dfrac{1(1)}{10}=\dfrac{1}{10}"

"E(X)=\dfrac{3}{10}(0)+\dfrac{6}{10}(1)+\dfrac{1}{10}(2)=\dfrac{8}{10}"

"E(X^2)=\dfrac{3}{10}(0)^2+\dfrac{6}{10}(1)^2+\dfrac{1}{10}(2)^2=\dfrac{10}{10}"

"Var(X)=E(X^2)-(E(X))^2=\dfrac{10}{10}-(\dfrac{8}{10})^2=0.36"


b.  With replacement.


"P(X=0)=\dfrac{3}{5}(\dfrac{3}{5})=\dfrac{9}{25}"

"P(X=1)=\dfrac{2}{5}(\dfrac{3}{5})+\dfrac{3}{5}(\dfrac{2}{5})=\dfrac{12}{25}"

"P(X=2)=\dfrac{2}{5}(\dfrac{2}{5})=\dfrac{4}{25}"

"E(X)=\dfrac{9}{25}(0)+\dfrac{12}{25}(1)+\dfrac{4}{25}(2)=\dfrac{20}{25}=\dfrac{4}{5}"

"E(X^2)=\dfrac{9}{25}(0)^2+\dfrac{12}{25}(1)^2+\dfrac{4}{25}(2)^2=\dfrac{28}{25}"

"Var(X)=E(X^2)-(E(X))^2=\dfrac{28}{25}-(\dfrac{4}{5})^2=0.48"